### v2

# M3Bench: Benchmarking Whole-body Motion Generation for Mobile Manipulation in 3D Scenes

Abstract
We propose M3Bench, a new benchmark of whole-body motion generation for mobile manipulation tasks. Given a 3D scene context, M3Bench requires an embodied agent to understand its configuration, environmental constraints and task objectives, then generate coordinated whole-body motion trajectories for object rearrangement tasks. M3Bench features 30k object rearrangement tasks across 119 diverse scenes, providing expert demonstrations generated by our newly developed M3BenchMaker. This automatic data generation tool produces coordinated whole-body motion trajectories from high-level task instructions, requiring only basic scene and robot information. Our benchmark incorporates various task splits to assess generalization across different dimensions and leverages realistic physics simulation for trajectory evaluation. Through extensive experimental analyses, we reveal that state-of-the-art models still struggle with coordinated base-arm motion while adhering to environment-context and task-specific constraints, highlighting the need to develop new models that address this gap. Through M3Bench, we aim to facilitate future robotics research towards more adaptive and capable mobile manipulation in diverse, real-world environments.
I Introduction
Humans possess an innate ability to manipulate their environment with remarkable flexibility and coordination, seamlessly integrating locomotion and manipulation. In contrast, robots still struggle to match this level of adaptability and proficiency in mobile manipulation tasks. Traditional trajectory optimization methods[1, 2], while capable of producing complex whole-body motion, are limited by their reliance on perfect environmental knowledge and predefined goal configurations (e.g., grasp poses), restricting their effectiveness and generalizability in real-world scenarios.
To address these limitations, researchers have turned to learning-based approaches. However, existing learning-based models deployed on mobile manipulators often address individual subproblems such as navigation, grasping, or motion planning in a modular fashion [3, 4]. These approaches can lead to suboptimal or infeasible solutions as they overlook the potential of coordinated whole-body motion and create misalignments between module outputs and requirements. For instance, in an object fetching task, a nearby navigable position might be impractical for the arm to reach the object, or a feasible grasp pose could be unachievable due to embodiment constraints (see Fig. 2(a)). These limitations underscore the need for a holistic approach that jointly understands robot embodiment, environment, and task objectives to generate feasible coordinated whole-body motion for mobile manipulators in diverse 3D scenes.
However, the development of such a holistic approach has been hindered by the scarcity of datasets capturing whole-body motions in 3D scenes. Existing benchmarks primarily focus on navigation [5] or task planning without actual object manipulation [6], lacking necessary data for studying embodied motion in 3D environments. Manipulation-oriented benchmarks often concentrate on object-centric modeling (e.g., grasping [7] and affordance [8, 9, 10]) with simplified embodiment, or stationary manipulation [11, 12] that neglects the coordination between mobile base and arm motions in diverse scenes.
To address this need, we developed M3BenchMaker (see Fig. 2(b)), an automatic data generation tool to produce whole-body motion trajectories as demonstrations for robot learning. M3BenchMaker procedurally generates coordinated trajectories from high-level task instructions, requiring only the action type, object link, and Unified Robot Description Format (URDF) of the scene and robot. It employs an energy-based model [13] to predict grasp pose or placement candidates and leverages an advanced virtual kinematics technique [1] to solve for coordinated whole-body motion trajectories (see Sec. III for details). This tool not only streamlines the demonstration collection process but also allows researchers to generate additional samples from customized robot and scene configuration for their own studies.
Leveraging M3BenchMaker, we create M3Bench, a comprehensive benchmark for task-oriented whole-body motion generation in household environments (see Fig. 1). M3Bench features challenging object rearrangement tasks that require a mobile manipulator to understand its embodiment, environmental contexts, and task objectives from 3D scans to produce coordinated motion for picking or placing objects across diverse scenes. It comprises 30k object rearrangement tasks involving 32 distinct object types across 119 diverse household scenes, covering a wide spectrum of environmental and task objective constraints on robot embodied motion. The benchmark provides expert demonstrations for each task generated by M3BenchMaker, along with additional metadata including natural language task instructions, panoptic maps, and ego-centric camera videos to facilitate research in related areas of embodied AI, such as embodied instruction following etc.
To facilitate in-depth evaluation of motion generation from 3D scans for mobile manipulation, M3Bench incorporates various task splits to assess generalization across different dimensions, such as novel scenes and objects. We leverage a realistic physics simulation platform [14] to evaluate the feasibility of generated motion trajectories and determine whether the robot can physically grasp the object or place it steadily at the desired location. Furthermore, our thorough experiments reveal that both hybrid approaches combining motion planning with affordance prediction and direct imitation learning methods struggle to effectively tackle mobile manipulation tasks across diverse 3D environments. This underscores the necessity of our benchmark for advancing research in whole-body motion generation for mobile manipulation in 3D scenes, paving the way for more adaptive and capable robotic systems in complex environments.
| Benchmark | Mobile Manipulation | Whole-body Demonstration | Procedural Generation | Household Scene | Language | Physical Grasp | Egocentric Perception | Flexible Material |
| ACRV [15] | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ | ✓ | ✗ |
| Alfred [16] | ✗ | ✗ | ✓ | 120 | ✓ | ✗ | ✓ | ✗ |
| ManiSkill [11, 17] | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | ✗ | |
| Calvin [18] | ✗ | ✗ | ✓ | ✗ | ✓ | ✓ | ✓ | ✗ |
| Behavior [19] | ✓ | ✗ | ✓ | 50 | ✗ | ✗ | ✓ | ✓ |
| RLBench [20] | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | |
| VLMbench [21] | ✗ | ✗ | ✓ | ✗ | ✓ | ✓ | ✗ | |
| Ravens [22] | ✗ | ✗ | ✓ | ✗ | ✓ | ✓ | ✗ | |
| MotionBenchMaker [23] | ✗ | ✗ | ✓ | ✗ | ✗ | ✗ | ✗ | |
| Habitat HAB [24] | ✓ | ✗ | ✓ | 105 | ✗ | ✗ | ✓ | ✗ |
| ARNOLD [25] | ✗ | ✗ | ✗ | 20 | ✓ | ✓ | ✓ | ✓ |
| Ours | ✓ | ✓ | ✓ | 119 | ✓ | ✓ | ✓ | ✓ |
In this work, we make three key contributions as follow:
We introduce, M3Bench, a comprehensive benchmark for task-oriented whole-body motion generation, facilitating the learning and evaluation of mobile manipulation in household environment.
We developed, M3BenchMaker, an automatic data generation tool that procedually generate whole-body motion trajectories from high-level task instructions, allowing users to easily generate data samples with customized robot and scene configurations.
We provided an in-depth evaluation of motion generation from 3D scans for mobile manipulation, revealing weaknesses of current arts in promoting future research in mobile manipulation across diverse 3D scenes.
These contributions collectively advance the state of the art in mobile manipulation, providing the tools and insights necessary to develop more capable and adaptive robotic systems for real-world applications.
II Related Work
Recent benchmarks for Embodied AI have focused on training agents for household activities [26, 27], often simplifying actions to symbolic operations [28] or navigation [5], which limits real-world applicability. While more recent benchmarks allow fixed-base manipulators to interact with objects via realistic simulation [17, 11, 20] or mobile agents to navigate and manipulate in diverse scenes [19, 24, 4], they lack expert data for whole-body motion in 3D scenes. Our M3Bench addresses this gap with a unique collection of whole-body motions for mobile manipulators in various 3D scenes, enabling models to comprehend environmental constraints and generate coordinated motions in complex environments. Table I compares M3Bench with other existing benchmarks.
In addition, unlike existing motion benchmark tools that focus on stationary manipulation [23] in simplified environments, our M3BenchMaker could procedually generate whole-body motion trajectories for mobile manipulators in diverse 3D scenes, facilitating practical model development for real-world tasks.
The proposed M3Bench and tool offer benefits for related fields in embodied AI. In visual affordance, recent studies focus on grasping poses [29, 7, 30] and functional parts [8, 9, 31, 32], but such object-centric modeling often lead to infeasible trajectories in complex scenes. Our whole-body motion data in 3D scenes enables modeling context-aware embodied visual affordance. Similarly, in embodied instruction following, existing benchmarks focus on navigation [5] or symbolic actions [16, 33]. Our M3BenchMaker allows researchers to generate whole-body motion data, facilitating models to produce low-level motion directly from language instructions.
III The M3BenchMaker
Diverse whole-body motion trajectories for mobile manipulators in complex 3D environments is crucial for advancing embodied AI. However, collecting expert demonstrations for training models are usually time-consuming and challenging. To address this, we introduce M3BenchMaker, a user-friendly tool that streamlines the generation of whole-body motion trajectories in 3D scenes, significantly reducing the time and effort required to create large-scale datasets for mobile manipulation tasks in various environments. Fig. 3 illustrates the architecture of the M3BenchMaker.
III-A Task Builder
The task builder serves as the primary user interface, allowing users to define manipulation tasks using high-level action commands such as picking, placing, and reaching. Users no longer need to manually specify grasping poses, placement locations, base positions, or create optimization programs for motion trajectories. To define a task, users simply select target object links from the scene URDF, set the robot’s initial position, and specify the desired action types. The task builder then creates an instance of the data generation pipeline, integrating subsequent modules to procedually generate whole-body motion trajectories. For enhanced dataset diversity, the task builder supports data augmentation via the Conditional Scene Sampler (see Sec. III-B). This feature facilitates the training and evaluation of embodied AI models in complex environments by generating varied scenarios from a single task definition.
III-B Conditional Scene Sampler
The conditional scene sampler generates diverse initial configurations for data augmentation by randomizing object and robot positions and orientations. It produces variations dependent on the original scene’s object relations, ensuring physical feasibility and contextual consistency required by the task. For instance, in a task involving picking an object from a table, the sampler ensures the sampled objects remains on table (see Fig. 3 orange box). This is achieved by recognizing supporting planes for objects and the robot through analysis of surrounding geometries.
To identify supporting planes, we parameterize a surface plane as , where is the normal vector, is the distance to origin, and defines the plane’s polygon outline. The most likely supporting plane for a bottom surface is identified by solving:
| s.t. | (2) | |||
where is a set of supporting plane candidates, denotes polygon area, computes intersection, and projects bottom surface points onto the supporting plane, and are distance and angle thresholds. Eq. 1 defines the contact ratio, while Eqs. 2 and 3 enforce alignment and distance constraints. We utilize the method in [34] to extract surface planes and solve the optimization problem by iteratively identifying the plane that maximizes Eq. 1 while satisfying the constraints.
III-C Goal Configuration Generator
This module efficiently generates 6D end-effector poses for grasping or placing target objects, serving as optimization objectives for motion planning. We employ an energy-based model to predict candidate goal configurations based on target object geometry [13]. However, this object-centric approach, which considers only object geometry without accounting for the robot’s kinematic constraints or environmental contexts, results in only a small subset of candidates being feasible for the task. To address the computational expense of evaluating all candidates through motion planning, we developed an adaptive sampling algorithm that efficiently draws samples from the candidate set, significantly accelerating the motion generation process.
Detailed in Alg. 1, our algorithm iteratively selects and updates the sampling probability of candidates based on their feasibility scores. It utilizes a K-D tree for efficient neighbor searching and initializes feasibility scores using the candidates’ energy values. By concentrating sampling in promising regions of the goal configuration space while maintaining exploration, the algorithm significantly reduces the number of expensive feasibility checks required to identify viable goal configurations.
III-D VKC Problem Generator
The VKC problem generator automates the construction of motion planning programs, formulating comprehensive optimization problems that encapsulate all necessary constraints and objectives for computing whole-body motion trajectories, utilizing task specifications, URDF, and goal configurations from preceding modules. We employ the VKC approach [1] to solve for whole-body motion of mobile manipulators, modeling the mobile base, robot arm, and manipulated object as a unified system, achieving superior foot-arm coordination through simultaneous optimization and surpassing traditional methods that separate base and arm planning.
Our implementation follows TrajOpt and ROS-Industrial Tesseract conventions [35], effectively incorporating kinematic constraints while avoiding large-space searches. The trajectory optimization minimizes joint travel distances and overall smoothness, with inequality constraints for joint limits, collision avoidance, and end-effector pose reaching. We adopt a sequential convex optimization method [36] to solve the resulting problem, yielding feasible, coordinated whole-body motion trajectories for diverse mobile manipulation tasks without manual task-specific planner programming.
By automating these processes, M3BenchMaker empowers researchers to efficiently collect tailored whole-body motion trajectories, significantly advancing embodied AI in complex 3D environments.
IV The M3Bench
The M3Bench aims to advance robot capabilities in coordinating whole-body movements within complex environments, inspired by human ability to seamlessly perform such tasks. It challenges mobile manipulators to generate coordinated whole-body motion trajectories for picking or placing everyday objects in 3D scenes, requiring agents to jointly understand their embodiment, environmental contexts, and task objectives from 3D scans.
IV-A Simulation Environment
Simulation Platform. Our benchmark, built on Isaac Sim [14], provides a high-fidelity physics simulation that meticulously models real-world properties and interactions. This platform enables precise evaluation of motion trajectory feasibility, grasping abilities, and the complex interplay between mobility and manipulation. Additionally, it could generate rich perceptual data (e.g., RGB-D image) that closely mimics the sensory input available to real-world robots.
Scene and Robot Configuration. The benchmark comprises 119 diverse household scenes containing 32 types of objects, curated from PhyScene [37]. These interactive 3D scenes are enhanced with physical properties and rich materials for photo-realistic and physics-realistic simulation. For the robot, we employ a common mobile manipulator configuration: a 7-DoF Kinova Gen3 robotic arm with a parallel gripper, mounted on an omnidirectional mobile base. This setup facilitates complex manipulations requiring coordinated base and arm movements.
IV-B Task Design and Variations
M3Bench focuses on two primary object rearrangement tasks: picking and placing. Given a 3D point cloud of the scene, a mask of the target object, and its initial configuration, the robot must generate whole-body motion trajectories to manipulate the object. The tasks are defined as: (i) Pick tasks: Navigate to, reach, and grasp a specified object from its initial location; (ii) Place tasks: Transport a held object to a designated location and place it stably. Success in both tasks requires avoiding collisions along the trajectory and maintaining the desired goal state for 2 seconds.
The task pool encompasses a wide range of mobile manipulation scenarios, featuring 32 object types with varying properties across 119 diverse household scenes. Each scene presents unique layouts, furniture arrangements, and obstacle configurations. Tasks are generated by selecting appropriate objects and placement locations based on scene categories. We employ the conditional scene sampler (Sec. III-B) to generate various initial configurations, further challenging the robot to generate coordinated whole-body motions while adapting to environmental constraints and task objectives.
IV-C Data Collection
Demonstration Generation. We utilized our developed M3BenchMaker to generate demonstrations for each task. The tool takes as input the scene and robot URDF, target object link, and task type (pick or place), then generates a whole-body motion trajectory for the robot. The optimization program in M3BenchMaker ensures these trajectories are collision-free and kinematically feasible. Each trajectory is then verified for physical feasibility in Isaac Sim, with only valid demonstrations and their corresponding tasks included in the benchmark. In total, we collected 30k valid demonstrations, each containing 30 waypoints.
Additional Metadata. To facilitate embodied AI research, we provide comprehensive metadata for each task (see Fig. 4). This includes annotations for all links in the scene URDF, covering object categories and simulation properties. We employ a template-based approach with lexicalized phrase candidates to generate language instructions for each task. For example, the template “Pick [object] in [room] on [position]” might be realized as “Pick the cup in the living room on the dining table”. During task execution, Isaac Sim’s built-in rendering capabilities, combined with annotated information, generate pixel-accurate semantic and instance segmentations along with egocentric camera views. This rich combination of annotations, trajectory data, and language instructions creates a comprehensive resource for exploring various aspects of embodied intelligence.
IV-D Benchmark
| Split | Pick | Place |
| Train | 14,793 | 7,478 |
| Val | 948 | 479 |
| Test | 3,225 | 1,630 |
| Novel Object | 688 | 397 |
| Novel Scene | 762 | 369 |
| Novel Scenario | 204 | 77 |
| Total | 20,620 | 10,430 |
| Statistics | Value |
| Bathroom | 132 |
| Bedroom | 198 |
| Kitchen | 97 |
| Living room | 129 |
| Total scenes | 119 |
| Object types | 32 |
| Total objects | 588 |
Data Split and Statistics. The tasks in M3Bench are carefully divided into several splits to assess different aspects of generalization capabilities. Objects and scenes are randomly categorized into seen and unseen subsets. The primary evaluation set, the Base split, encompasses all seen objects and scenes, divided into Train (75%), Val (5%), and Test (20%) sets. Three additional splits challenge model generalization: Novel Object (unseen objects in seen scenes), Novel Scene (seen objects in unseen scenes), and Novel Scenario (unseen objects in unseen scenes). Tables III and III present detailed statistics of these splits and task configurations, enabling systematic evaluation of model generalization across various dimensions of mobile manipulation in 3D scenes.
Metrics. We employ a multi-faceted approach to evaluate motion generation models. Task success rate serves as the primary metric, determined by the robot’s ability to complete specified tasks and maintain the desired state for 2 seconds, as verified by the Isaac Sim physics engine. We also measure the closest distance from the end-effector to the target as an auxiliary metric, reflecting the trajectory’s effectiveness in reaching the object or placement location. To assess trajectory quality, we utilize several quantitative measures: environment collision, self-collision, joint limit violation, and trajectory solving time. This comprehensive set of metrics evaluates models’ capabilities in generating effective and efficient motion trajectories for mobile manipulation in 3D scenes.
V Experiments
V-A Experimental Setup
Models for M3Bench. Due to the lack of existing models for whole-body motion generation in mobile manipulation within 3D scenes, we adapt three state-of-the-art approaches to our benchmark:
Implementation Details. For MNet and MFormer, we generate 3D scans from scene URDF. To enhance learning tractability, we apply a perception bounding box around the robot and target object to crop the scans, focusing the model’s attention on relevant spatial information. We train MNet and MFormer on the Train split and perform model selection on Val. In constrast, as ModularMP does not involve learning procedure, we evaluate it directly on the Test and Novel splits. To simplify the optimization problem in ModularMP, we ignore collisions between the end-effector and target object during motion planning, as considering these collisions would frequently result in infeasible trajectories.
V-B Experimental Results
| Test Split | Method | Pick Task | Place Task | ||||||||||
| Succ(%) | Dist(m) | JVio(%) | EnvColl(%) | SelfColl(%) | Time(s) | Succ(%) | Dist(m) | JVio(%) | EnvColl(%) | SelfColl(%) | Time(s) | ||
| Base Test | MNet | ||||||||||||
| MFormer | |||||||||||||
| ModularMP | |||||||||||||
| Novel Object | MNet | ||||||||||||
| MFormer | |||||||||||||
| ModularMP | |||||||||||||
| Novel Scene | MNet | ||||||||||||
| MFormer | |||||||||||||
| ModularMP | |||||||||||||
| Novel Scenario | MNet | ||||||||||||
| MFormer | |||||||||||||
| ModularMP | |||||||||||||
The experimental results are summarized in Tab. IV. Trajectories are evaluated in Isaac Sim using metrics described in Tab. III. Particularly, for the ModularMP model, when motion planning fails to solve the problem (i.e., optimization does not converge), we consider it as a failure instance.
Across Models. Comparing the baseline models, we observed that ModularMP outperforms other baselines in both pick and place tasks, demonstrating higher success rates and lower distances to goal. This aligns with our hypothesis that integrating conventional motion planning with affordance prediction would better generalize across diverse 3D scenes. However, ModularMP’s superior performance comes at the cost of significantly increased computation time, primarily due to optimization complexity in large-scale 3D environments. Additionally, its effectiveness is constrained by the quality of predicted grasp and placement poses, as inappropriate predictions may lead to optimization failures or environmental collisions (see Fig. 2(a)). Despite ModularMP’s relative success, the overall low success rates indicate that combining conventional motion planning with affordance prediction is insufficient.
On the other hand, learning-based models MNet and MFormer, while more time-efficient, struggle to produce feasible whole-body trajectories in complex scenes. Their performance is particularly poor, with near-zero success rates even in the Base split containing familiar objects and scenes. The generated trajectories often violate joint limitations and cause collisions, indicating a fundamental inability to adapt to complex 3D environments and produce feasible whole-body motions. These findings highlight the persistent challenge of generating whole-body motion trajectories for robots in complex 3D scenes. Further research is needed to enhance robot capabilities in manipulating complex environments.
Across Tasks. The experiment results reveal distinct performance patterns between pick and place tasks. While ModularMP maintains better performance in both tasks, its success rates significantly drop in place tasks, and all models require more time to generate trajectories for the place tasks. This discrepancy suggests that generating coordinated whole-body motion trajectories for placing objects is more challenging than for picking objects, as it involves additional constraints such as stable placement locations, appropriate object orientation, and reachable motion trajectory. The increased complexity of place tasks explains the lower success rates and longer execution times observed across all models.
On Generalization. While the learning-based methods MNet and MFormer exhibit near-zero success rates across all test splits, including the Base and generalization splits, we observe significantly higher distances from the end-effector to goal in the Novel splits, indicating these models’ struggle to adapt to unseen objects and scenes. Particularly, the distance to goal in unfamiliar scenes (Novel Scene and Novel Scenario splits) exceeds that of the Novel Object split, suggesting that the impact of novel scenes is more significant than novel objects. In contrast, the conventional motion planning method ModularMP demonstrates consistent performance across all splits. Nevertheless, their low success rates across splits underscore the considerable difficulties of mobile manipulation in diverse household environments.
Remarks. Our experimental results reveal two crucial insights:
While combining motion planning with affordance prediction shows consistent performance across all splits, its overall success rates remain low. This indicates that even advanced hybrid approaches fail to fully address the challenges of mobile manipulation in diverse 3D scenes, underscoring the urgent need for more sophisticated models capable of tackling complex mobile manipulation tasks.
Current arts of learning methods that directly imitate whole-body trajectories fail dramatically when applied to mobile manipulation across large-scale 3D scenes and objects. Their near-zero success rates across all splits highlight the necessity for significant advancements in two key areas: (a) the development of increasingly fine-grained representations for perceptual inputs to better capture the complexity of 3D environments, and (b) more sophisticated modeling of continuous whole-body motions to generate feasible trajectories in challenging scenarios.
VI Conclusion
We introduced M3Bench, a comprehensive benchmark for whole-body motion generation in mobile manipulation tasks across diverse 3D environments, featuring 30k object rearrangement tasks in 566 household scenes. We developed M3BenchMakerto efficiently generate whole-body motion trajectories from high-level instructions. Our experiments revealed significant limitations in current approaches, with even the best-performing hybrid method achieving low success rates and learning-based methods struggling dramatically. While M3Benchprovides a solid foundation for future research, several challenges remain: bridging the gap to real-world scenarios, extending to long-horizon tasks, expanding language instruction diversity, developing more sophisticated models for 3D context understanding and coordinated motion generation, facilitating research into versatile agents, and scaling up demonstrations. Addressing these challenges will be crucial for advancing the field towards more capable and adaptive robotic systems in diverse, complex, unstructured environments.
Acknowledgement: This work was supported in part by the National Natural Science Foundation of China (No. 62403064, 62403063). We thank the all the colleagues in Robotics Lab from BIGAI for fruitful discussions.